OpenStudy (anonymous):

state the domain and range f(x)=-2e^2x

myininaya (myininaya):

domain is all real numbers. now if we find the domain of the inverse (since f is 1 to 1) we can find the range of f(x) -2/y=e^(2x) ln(-2/y)=2x so x=1/2*ln(-2/y) f^-1(x)=1/2*ln(-2/x) the domain of f' is x>0 so the range of f(x) is [-inf,0)